### v2

@@ 4.4 Model Ensemble for Clean and Adversarial Image Classification
9:     Ensemble the defense model $f_{i}$ and the target model $f_{t}$ using Eq. 15.
10:  end for
5 Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies on two datasets and compare CAD to baselines from three research streams. For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy against each attack after adaptation for defense performance evaluation. (2) Average accuracy of all attacks occurring within and before each stage, i.e. average incremental accuracy [27], for assessing the knowledge retention of the model.
Datasets
CIFAR-10 [17] is a widely used dataset for adversarial attack and defense, containing 5000 images for training and 1000 images for testing in each of the 10 classes. ImageNet-100 [33] is a subset of ImageNet [4] and contains 100 classes with 1000 images for training and 100 images for testing in each class. Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively. Before training, the above datasets endured data augmentation including horizontal flip and random crop.
Attack Algorithms
We choose PGD-$l_{\infty}$ [23] as the initial attack. After that, we select 8 adversarial attack methods under $l_{\infty}$: BIM [18], RFGSM [34], MIM [6], DIM [40], NIM [21], SNIM [21], VNIM [38] and VMIM [38] to compose the attack pool in defense performance evaluation. The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [44]. As for evaluating the capability of CAD in migrating catastrophic forgetting, $l_{\infty}$ attack VMIM [38] and SNIM [21], $l_{2}$ attack CW [1] and DeepFool [25], $l_{1}$ attack EAD [2] and EADEN [2], and $l_{0}$ attack OnePixel [31] and SparseFool [24] compose the attack pool. Details of attack parameters are given in the supplementary.
Stage 0 1 2 3 4 5 6 7 8 -
Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
None-defense 0.6 1.3 0.0 0.0 0.0 0.0 0.2 0.0 0.0 96.4
TRADES [44] 55.8 60.2 57.4 55.9 55.8 57.1 57.1 58.9 57.3 84.8
JEM [10] 55.8 62.0 46.5 52.3 52.4 51.1 51.4 57.0 48.2 91.7
PAT [19] 45.5 45.1 43.8 45.5 45.3 45.2 45.2 43.9 45.4 69.3
GAIRAT [45] 66.9 70.4 66.4 66.9 66.7 66.5 66.5 68.6 66.4 89.4
FastAdv [14] 37.1 39.6 36.8 37.0 37.7 37.7 54.6 36.4 37.6 75.2
RPF [5] 60.9 52.8 52.2 58.9 54.9 57.6 57.5 53.5 52.6 83.5
DMAT [39] 71.7 76.0 73.1 71.7 71.7 73.0 73.0 75.5 73.0 92.4
EBM [12] 75.2 72.9 73.0 76.6 76.5 72.4 73.9 74.0 73.8 86.8
ADP [41] 85.4 80.5 79.3 84.4 85.1 84.1 83.7 81.0 79.1 80.5
DiffPure [26] 72.5 70.2 69.8 70.1 72.2 73.4 73.5 72.7 69.4 89.3
CAD(ours) 94.5 95.4 93.4 94.7 95.0 93.7 93.7 95.2 93.7 95.8
Table 1: Classification accuracy(%) against various attacks($l_{\infty}$) on CIFAR-10. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by efficient memory and few-shot defense feedback.
Implementation Details
Both the target model and defense model use ResNet as the backbone network in our method. Specifically, we use WideResNet-28-10 [42] for CIFAR-10 following [10] and ResNet-50 [11] for ImageNet-100 following [19]. A 4-layer ConvNet is used as the weight estimator model. The cosine classifier [8] is adopted as the classifier of the defense model. We utilize Torchattacks [15] for generating adversarial images. The amount of feedback per class is set to $K=10$. The number of stages is set to $T=8$ since there are 9 attacks (PGD is the initial attack). Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following [47]. Epochs for training $f_{0}$ are set to $100$ and fine-tuning epochs are set to $4$.
Defense Baselines
We first compare to adversarial training methods TRADES [44], JEM [10], GAIRAT [45], FastAdv [14], RPF [5] and DMAT [39], and purification methods EBM [12], ADP [41] and DiffPure [26] in defense performance evaluation. As for evaluating the capability of CAD in migrating catastrophic forgetting, we compare to CL methods EWC [16], SSRE [51], BEEF [36], and FACT [47]. Besides, we propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback. In premium-CAD, we train additional attack-specific defense models using abundant feedback in each stage, and ensemble them with the target model by the weight estimator. Details about the implementation of defense baseline and premium-CAD are given in the supplementary.
5.1 Comparisons to Baselines
To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks. For CAD, the defense environment is dynamic, with attacks occurring at different stages. The other baseline methods operate in a static defense environment that does not differentiate the order of attacks. All the evaluation is conducted on the whole test set of CIFAR-10 and ImageNet-100. Classification accuracy is reported in Tab. 1 and Tab. 2.
Stage 0 1 2 3 4 5 6 7 8 -
Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
None-defense 0.2 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 89.8
TRADES [44] 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
GAIRAT [45] 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
RPF [5] 56.4 49.9 51.3 56.6 54.4 52.6 53.9 53.3 50.1 69.4
DMAT [39] 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
ADP [41] 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
DiffPure [26] 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
CAD(ours) 83.7 83.4 82.0 82.5 83.5 82.0 81.9 82.2 81.9 88.1
Table 2: Classification accuracy(%) against various attacks($l_{\infty}$) on ImageNet-100.
Stage 0 1 2 3 4 5 6 7 8
Attack PGD SNIM VNIM CW DeepF. EAD EADEN OnePixel SparseF.
None-defense 0.6 1.3 0.0 0.0 8.6 0.0 0.2 39.1 29.7
EWC [16] 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
SSRE [51] 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
BEEF [36] 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
FACT [47] 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
CAD(ours) 94.5 89.1 90.5 84.1 84.8 81.1 82.1 82.8 78.2
Table 3: Continual defense performance (Average incremental accuracy(%)) against attacks under different norms $l_{p=0,1,2,\infty}$ on CIFAR-10. DeepF. and SparseF. denote DeepFool [25] and SparseFool [24] respectively.
As depicted in Tab. 1 and Tab. 2, despite the small scale of feedback, both CAD and premium-CAD demonstrate strong adaptability to each type of attack. When compared with the top baseline methods DMAT [39] and ADP [41], CAD outperforms them by approximately $10\%$ on CIFAR-10 and $6\%$ on ImageNet-100 for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $95.8\%$ accuracy on CIFAR-10 and $88.1\%$ on ImageNet-100, while the target model stands at $96.4\%$ on CIFAR-10 and $89.8\%$ on ImageNet-100. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
To assess the effectiveness of CAD in retaining knowledge, we compare it to CL baseline methods through experiments on CIFAR-10 with attacks under different norms $l_{p=0,1,2,\infty}$. All baseline methods are functioned within our dynamic scenario. As depicted in Tab. 3, both CAD and premium-CAD maintain robust defense performance despite changes in attack types. However, as the variety of attack types increases, the defense performance gradually diminishes. This indicates that defending against an expanding array of attack types presents a more challenging task for continual defense. Compared with the top baseline method FACT [47], CAD consistently outperforms it by an average of approximately $3.4\%$, highlighting CAD’s effectiveness in mitigating catastrophic forgetting.
It is noteworthy that premium-CAD surpasses all baseline methods. Given sufficient defense feedback and memory resources, premium-CAD emerges as the superior choice.
(a) PGD
(b) CW
(c) EAD
(d) OnePixel
Figure 3: T-SNE diagrams of features extracted from adversarial test data encoded by the defense model after all stages. Each color represents a class of CIFAR10. After several steps of adaptation, CAD retains the capability to categorize the attacked data.
5.2 Ablation Study
As argued, embedding reservation facilitates easier adaptation of the defense model and mitigates overfitting. Prototype augmentation preserves the decision boundary of previous stages, while model ensemble ensures robust classification performance for both clean and adversarial images. Therefore, we proceed to conduct an ablation study using 4 attacks PGD($l_{\infty}$), CW($l_{2}$), EAD($l_{1}$), and
Stage 0 1 2 3 -
Attack PGD CW EAD OnePixel Clean
w/o $L_{\mathrm{virtual}}$ 94.5 82.4 76.1 71.8 95.8
w/o $L_{\mathrm{proto}}$ 94.4 89.5 80.0 78.5 95.8
w/o $f_{p}$ 94.5 89.9 86.6 85.3 55.9
CAD(ours) 94.5 89.9 86.6 85.3 95.8
Table 4: Average incremental accuracy(%) on CIFAR-10 for ablation.
OnePixel($l_{0}$) to validate the efficacy of each component. As shown in Tab. 4, CAD without $L_{\mathrm{virtual}}$ and $L_{\mathrm{proto}}$ experiences a decrease of $13.5\%$ and $6.8\%$ respectivily in the last stage. The absence of model ensemble leads to the inability to uphold performance on clean images.
Additionally, we provide T-SNE diagrams of features extracted from adversarial test data in Fig. 3 to visualize the defense capability of CAD after adaptation. These features are encoded by the defense model after the last stage of adapting to OnePixel attack. From (a) to (d) in Fig. 3, the clustering of features becomes less and less discernible, indicating that the difficulty of few-shot adaptation increases with the introduction of new attack types. After adaptation, CAD maintains its defense performance against previous attacks, illustrating its ability to mitigate catastrophic forgetting as demanded in Principle 1.
5.3 Multiple Attacks in One Stage
In practice, it is common for multiple attacks to occur simultaneously within a
Stage 0 1 2 3
$\mathrm{number}=1$ 94.5 95.4 93.4 94.7
$\mathrm{number}=2$ 94.5 94.4 93.5 92.8
$\mathrm{number}=3$ 94.5 93.9 93.6 92.5
Table 5: Classification accuracy(%) about the number of attacks($l_{\infty}$) in one stage.
single stage. To assess CAD’s adaptability in such scenarios, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks on CIFAR-10. As presented in Tab. 5, when the number of attacks increases to $2$ and $3$, the classification accuracy against these attacks still remains at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively. The results indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
5.4 Number of Shots for Adaptation
As indicated in Principle 2, the defense adaptation relies on few-shot defense feedback. To investigate the influence of the amount of feedback $K$ per class, we
Stage 0 1 2 3
Attack PGD CW EAD OnePixel
$K=15$ 94.5 90.0 88.4 85.5
$K=10$ 94.5 89.9 86.6 85.3
$K=5$ 94.5 89.4 86.1 84.6
$K=1$ 94.5 87.9 84.6 83.4
Table 6: Average incremental accuracy(%) with different feedback scale.
conducted tests with $K=15,10,5,1$ on CIFAR-10. The results presented in Tab. 6 illustrate that defense performance diminishes as $K$ decreases. Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($93.4\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
5.5 Memory Allowance for Adaptation
Cache Size Model Size
Stage 1 2 3 1 2 3
CAD 50.3K 75.5K 100.4K 141M 141M 141M
$\mathrm{\textbf{CAD}}^{\dagger}$ 257M 377M 504M 280M 419M 558M
stage on CIFAR-10. As depicted in Tab. 7, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.
6 Conclusion
The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks. In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework. CAD is engineered to collect few-shot defense feedback from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles including few-shot feedback and memory-efficient adaptation, and CAD demonstrates strong performance when operating under these principles. Through experiments conducted on CIFAR-10 and ImageNet-100, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
A Weight Estimator Base on Self-Perturbation
In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation [37] to train a weight estimator to ensemble the target model and the defense model. We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples. Then, we train a 4-layer ConvNet as the weight estimator $f_{p}$ in a binary classification manner. In this section, we introduce the details of self-perturbation and model ensemble.
@@ A Weight Estimator Base on Self-Perturbation
B Premium-CAD
@@
4:     Ensemble the defense models $[f_{0},f_{1},\cdots,f_{i}]$ and the target model $f_{t}$ using Eq. 15.
@@
C Implementation Details
C.1 Adversarial Attack
In this work, all of the adversarial attacks are implemented using TorchAttacks [15]. The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [44]. Iterative steps and step size of iterative attacks are set to $10$ and $2/255$ respectively. The number of perturbed pixels of OnePixel [31] is set to $50$. The other parameters are the default ones in TorchAttacks.
C.2 Defense Baselines
The adversarial training baselines are implemented using RobustBench [3], while purification baselines are implemented using open-source code and model checkpoints. CL baselines are implemented using PyCIL [48]. The backbone models of both adversarial training methods and CL methods are standardized as WideResNet-28-10 for CIFAR-10 and ResNet-50 for ImageNet-100. CL models are trained under the same settings as CAD. Specifically, the training epochs and the amount of feedback per class for training CL models are set to 100 and K=10, respectively.

### v3

@@ Conclusion
9:     Refine the classifier using cached prototypes $P_{e}$.
10:     Ensemble the defense model $f_{i}$ and the target model $f_{t}$ using Eq.9.
11:  end for
Appendix A More Details about CAD
The overall algorithm of the CAD framework is presented in Algorithm 1.
Model Ensemble Base on Self-Perturbation
In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation (Wang et al. 2023a) to train a weight estimator to ensemble the target model and the defense model. We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples. Then, we train a 4-layer ConvNet as the weight estimator $f_{p}$ in a binary classification manner. In this section, we introduce the details of self-perturbation and model ensemble.
@@ Appendix A More Details about CAD
Stage 0 1 2 3 4 5 6 7 8 -
Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
None-defense 0.6 1.3 0.0 0.0 0.0 0.0 0.2 0.0 0.0 96.4
TRADES (Zhang et al. 2019) 55.8 60.2 57.4 55.9 55.8 57.1 57.1 58.9 57.3 84.8
JEM (Grathwohl et al. 2020) 55.8 62.0 46.5 52.3 52.4 51.1 51.4 57.0 48.2 91.7
PAT (Laidlaw, Singla, and Feizi 2020) 45.5 45.1 43.8 45.5 45.3 45.2 45.2 43.9 45.4 69.3
GAIRAT (Zhang et al. 2020) 66.9 70.4 66.4 66.9 66.7 66.5 66.5 68.6 66.4 89.4
FastAdv (Jiang et al. 2023) 37.1 39.6 36.8 37.0 37.7 37.7 54.6 36.4 37.6 75.2
RPF (Dong and Xu 2023) 60.9 52.8 52.2 58.9 54.9 57.6 57.5 53.5 52.6 83.5
DMAT (Wang et al. 2023b) 71.7 76.0 73.1 71.7 71.7 73.0 73.0 75.5 73.0 92.4
EBM (Hill, Mitchell, and Zhu 2020) 75.2 72.9 73.0 76.6 76.5 72.4 73.9 74.0 73.8 86.8
ADP (Yoon, Hwang, and Lee 2021) 85.4 80.5 79.3 84.4 85.1 84.1 83.7 81.0 79.1 80.5
DiffPure (Nie et al. 2022) 72.5 70.2 69.8 70.1 72.2 73.4 73.5 72.7 69.4 89.3
CAD(ours) 94.5 95.4 93.4 94.7 95.0 93.7 93.7 95.2 93.7 95.8
Table 7: Classification accuracy(%) against various attacks($l_{\infty}$) on CIFAR-10. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by efficient memory and few-shot defense budget.
Appendix B Premium-CAD
@@
4:     Ensemble the defense models $[f_{0},f_{1},\cdots,f_{i}]$ and the target model $f_{t}$ using Equation 14.
@@
Stage 0 1 2 3 4 5 6 7 8
Attack VMIM VNIM NIM DIM MIM RFGSM BIM SNIM PGD
CIFAR 94.8 94.6 93.4 93.5 94.2 93.4 92.5 92.2 93.3
ImageNet 82.5 82.1 82.0 82.0 92.4 82.2 81.9 82.3 81.7
Table 8: Classification Acc (%) against changed attack sequence.
@@
Appendix C More Implementation Details
Adversarial Attack
In this work, all adversarial attacks are implemented using TorchAttacks (Kim 2020). The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following (Zhang et al. 2019). Iterative steps and step size of iterative attacks are set to $10$ and $2/255$ respectively. The number of perturbed pixels of OnePixel (Su, Vargas, and Sakurai 2019) is set to $50$. The other parameters are the default ones in TorchAttacks.
Defense Baselines
The adversarial training baselines are implemented using RobustBench (Croce et al. 2021), while purification baselines are implemented using open-source code and model checkpoints. CL baselines are implemented using PyCIL (Zhou et al. 2023). The backbone models of both adversarial training methods and CL methods are standardized as WideResNet-28-10 for CIFAR-10 and ResNet-50 for ImageNet-100. CL models are trained in the same settings as in CAD. Specifically, the training epochs and the amount of feedback per class for training CL models are set to 100 and K=10, respectively.
Appendix D More Experimental Results
Continual Adversarial Defense on CIFAR-10
We compare our CAD with more adversarial defense baselines on CIFAR-10. As presented in Table 7, both CAD and premium-CAD demonstrate strong adaptability to each type of attack. When compared with the top baseline methods DMAT (Wang et al. 2023b) and ADP (Yoon, Hwang, and Lee 2021), CAD outperforms them by approximately $10\%$ for each attack. As for clean images, CAD achieves $95.8\%$ accuracy while the target model stands at $96.4\%$.
Change the sequential of attacks
For different attack sequences, we reverse the sequential of attacks and conduct an experiment. As shown in Table 8, when the attack sequence changes, CAD maintains a consistent defense performance, indicating our CAD is agnostic to the attack order.
Change the intensity of attacks
Stage 0 1 2 3 4 5 6
ImageNet-$\epsilon/255$ $4$ $6$ $8$ $10$ $12$ $14$ $16$
Acc 83.7 83.6 83.3 83.3 83.4 82.2 82.5
CIFAR-$\epsilon/255$ $8$ $10$ $12$ $14$ $16$ $18$ $20$
Acc 94.5 94.4 94.4 93.9 94.0 93.2 93.4
Table 9: Classification Acc (%) against various attack intensity of PGD($l_{\infty}$).
For different attack intensities, we conduct experiments using PGD($l_{\infty}$) with various $\epsilon$. As shown in Table 9, when the attack intensity changes, CAD maintains a consistent defense performance, indicating our CAD is agnostic to the attack intensity.
Averaged accuracy when the number of classes increases
We present curves depicting the relationship between averaged accuracy and the number of classes in Figure 4. As the number of classes increases, CAD continues to perform well on both the new attack and previous attacks.
Figure 4: Average Acc with 10 classes and 100 classes per stage for CIFAR10 and ImageNet100.
Appendix E Limitations and Future Work
The increase in attacks denotes the expansion of the defense model classifier, especially when tackling large datasets such as the original ImageNet. Although CAD refines the defense budget before each stage’s classifier expansion, too many classes can lead to an undesirable decrease in classification performance. In the future, we will focus on this issue and refine the classifier to guarantee the consistent performance of the defense model.